# Inertial Confinement Fusion Forecasting via LLMs

### v2

@@ 3.1 Main Results
  # of samples Ours HoGRC [43] NGRC [20] Time-LLM [31]
@@
  # of epochs Ours HoGRC [43] NGRC [20] Time-LLM [31]
@@
3.2 Diagnostic Experiment
This section ablates Fusion-LLM’s systemic design on val split of Fusion4AI. All experiments use the Llama 3 8B variant. Appendix §S2 has more experimental results.
Key Component Analysis. We first investigate the two principal modules of Fusion-LLM, specifically, Fusion-Specific Prompt and Signal-Digesting Channels. We construct a baseline model with generic dummy prompts which only provide broad, non-specific instructions regarding fusion, and a rudimentary encoder composed of a single linear layer. As shown in Table 2a, the baseline model achieves 3.57 CAE. Upon applying Fusion-Specific Prompt to the baseline, we observe significant improvements for CAE from 3.57 to 2.59. Furthermore, after incorporating Signal-Digesting Channels into the baseline model, we achieve significant gains of 1.56 CAE. Finally, by integrating both core techniques, our Fusion-LLM delivers the best performance of 1.19 CAE. These findings affirm that the proposed Fusion-Specific Prompt and Signal-Digesting Channels operate synergistically, and validate the effectiveness of our comprehensive model design.
@@ 3.2 Diagnostic Experiment
Reservoir and LLM Comparison. To thoroughly explore the training effectiveness of our methodology under conditions of limited sample availability, we perform a comparative analysis using a variable number of training samples with two concurrent reservoir methods [43, 20] and one LLM-based method [31] in Table 2d using CAE. The empirical findings demonstrate that our approach consistently outperforms all competing strategies across various sample configurations. Notably, this superior performance and training effectiveness are evident even with as few as 20 samples. Such robust efficacy is critical in scientific AI applications, where datasets are often constrained in size.
In addition, we study the training efficiency of our approach in contrast to the above strong baselines [43, 20, 31] in Table 2e, across various training epochs. The experimental outcomes illustrate that our approach not only outperforms its counterparts but also demonstrates superior efficiency. Specifically, our method is capable of achieving comparable or superior performance in a significantly reduced training duration. For instance, our model requires only 10 epochs to achieve better performance than other methods that require 100 epochs. This enhanced efficiency in training is particularly significant, as it demonstrates the potential of our approach to deliver robust performance swiftly, thereby facilitating more expedient research in practical scenarios.
@@ 4 Related Work and Discussion
AI ${}_{\!}$for ${}_{\!}$cience. ${}_{\!}$ ${}_{\!}$is ${}_{\!}$increasingly ${}_{\!}$acknowledged ${}_{\!}$as ${}_{\!}$a ${}_{\!}$critical ${}_{\!}$instrument and recently ${}_{\!}$led trending of ${}_{\!}$significant scientific discovery ${}_{\!}$[32, 40, 55, 5]. ${}_{\!}$he ${}_{\!}$trajectory ${}_{\!}$of ${}_{\!}$ in ${}_{\!}$scientific ${}_{\!}$research ${}_{\!}$commenced ${}_{\!}$with ${}_{\!}$different ${}_{\!}$elementary data analysis methods (e.g., Rule-based [11, 60], Bayesian [19], Analogy [24, 30, 76], Evolutionary [35, 16], Connectionism [83, 40], etc) and has evolved into the advanced foundation models [23, 81, 15, 18, 10]. This development is underscored by considerable progress in AI recently, particularly the emergence of sophisticated constructs such as LLMs [2, 78, 75, 95], which have fundamentally altered our approach to scientific challenges. By incorporating expansive knowledge bases, AI is indispensable for deciphering vast and varied scientific data.
The uniqueness of our work lies in the challenging nature of accessing ICF data, given the difficulty in observing LPI, which renders conventional AI tools less effective in this crucial area of physics. Building on the success of LLMs, this study harnesses their extensive knowledge base to unlock robust predictive capabilities for physical phenomena. Our empirical findings validate our methodological approach, positioning our work as pioneering and innovative. This not only demonstrates the applicability of LLMs in the scientific domain but also fosters a dynamic interplay between AI and scientific exploration. We showcase the potential of our LLM-based solution to make meaningful contributions in addressing complex scientific challenges, thus advancing both AI and science.
@@
Plasma Physics for Fusion. In the realm of plasma physics, particularly in ICF, efficient and symmetric driving of a target is imperative. Specifically, LPI represents significant impediments due to the complex dynamics of phenomena in fusion. For instance, Stimulated Raman and Brillouin Scattering (SRS and SBS) can cause the reflection of laser beams [37]. Moreover, Cross-Beam Energy Transfer [28] may adversely influence the symmetry of implosion . A critical issue particularly in direct drive scenarios is the preheating caused by hot electrons produced via SRS and Two-Plasmon Decay, which potentially increases the shell entropy and diminishes the implosion efficiency [68, 21, 14, 54]. Conversely, these hot electrons might contribute positively by depositing their energy in the compressed shell, thereby enhancing the ignition shock and aiding ignition processes such as those observed in shock ignition, a novel high-gain strategy [9, 53]. Understanding the dynamics ofLPI and establishing predictive models for hot electron generation in direct drive configurations are crucial.D In response to these issues and challenges, extensive physical experimentation and simulations [68, 21, 14, 54, 9, 53] are necessary, which are both costly and time-consuming.
To address the above challenges, our AI-based approach emerges as a potential alternative. By reformulating LLMs with a simple yet effective pipeline, our model can effectively engage in such ICF tasks as Computational Reservoir to encapsulate domain-specific knowledge and generalize within the domain. This capability potentially allows LLMs to guide ICF design, circumventing the need for extensive and expensive experimental setups and simulations. Empirical evidence indicates that the incorporation of LLMs had promising to revolutionize the predictive modeling in plasma physics by providing quick, cost-effective insights that are grounded in generalizable knowledge.
Reservior Computing. The ICF data exhibits a time-series nature, where laser intensity influences the target, correlating with the volume of hot electron energy emission. Traditional machine learning approaches [72, 7, 3, 84, 61, 62, 33] for time-series data typically involve dynamically transforming temporal inputs into a high-dimensional state space through nonlinear mappings. Revolutionary advancements in the field, such as echo state networks [29, 47] and liquid state machines [49, 92], have introduced the concept of Reservoir Computing (RC). RC presents an innovative framework [51, 63, 20, 48] characterized by a static “reservoir”. This reservoir is pre-trained, reducing the computational expenses associated with training only the readout parameters, rather than the entire network. This feature makes RC advantageous for applications involving temporal data [48, 82]. In this paradigm, the input signals are transformed non-linearly into a dynamic time-series system’s state space, from which outputs are linearly extracted [74].
In contrast to traditional reservoirs (e.g., RNN-based [20, 48]) that often struggle with processing highly dynamic temporal data, and contemporary time-series Large Language Models (LLMs) [79, 4, 31, 94] that employ massive parameters but focus solely on temporal sequences, often neglecting the vast knowledge embedded within LLMs, we propose a re-imagining of LLMs within the RC paradigm. This innovative integration not only leverages extensive domain-specific knowledge and adept handling of temporal data but also reduces the need for extensive meta-parameter tuning by maintaining most parameters as static during training. Consequently, our approach emerges as a compelling alternative, effectively addressing the challenges posed by both traditional RNN-based reservoirs and modern time-series LLMs, showcasing superior capability in handling intricate and large-scale temporal patterns.
@@ 5 Conclusion
Fusion energy stands as a pivotal pathway toward advancing human civilization to a Type I status on the Kardashev scale [34]. The key to realizing this potential lies in mastering Inertial Confinement Fusion, where understanding laser-plasma instabilities is paramount. To address this challenge, we present Fusion-LLM, a groundbreaking framework merging LLMs with reservoir computing. Our approach not only provides a cost-effective solution but also emerges as a top-tier contender in forecasting hot electron dynamics, offering invaluable insights for plasma scientists in ${}_{\!}$refining ${}_{\!}$ ${}_{\!}$designs. ${}_{\!}$eyond ${}_{\!}$its ${}_{\!}$immediate ${}_{\!}$impact ${}_{\!}$on ${}_{\!}$, ${}_{\!}$employing ${}_{\!}$s for ${}_{\!}$scientific ${}_{\!}$exploration ${}_{\!}$holds ${}_{\!}$promise ${}_{\!}$for ${}_{\!}$${}_{\!}$cross-domain ${}_{\!}$applications, ${}_{\!}$potentially ${}_{\!}$catalyzing ${}_{\!}$advancements ${}_{\!}$in ${}_{\!}$-driven ${}_{\!}$scientific ${}_{\!}$endeavors.
@@
The overall pipeline of Fusion-LLM is shown in Fig. 3. Experiments are conducted on two NVIDIA A100-40GB GPUs. For our approach, we keep all parameters of the LLMs and most of the SDC frozen during the fine-tuning. Only parameters pertaining to the Prediction Head and partial Spatial Encoder are trainable. The codes and dataset shall be publicly released upon paper acceptance.
  • Fusion-LLM is built from Llama-2-7B-hf / 3-8B [78, 4] to construct reservoir without tuning.
@@
  • Signal-digesting channels are composed of two components, temporal encoder and spatial encoder. The ${}_{\!}$former ${}_{\!}$one, ${}_{\!}$which incorporates 24 Transformer layers and a linear layer, captures temporal features ${}_{\!}$over ${}_{\!}$the ${}_{\!}$input ${}_{\!}$laser ${}_{\!}$signal. ${}_{\!}$his ${}_{\!}$module ${}_{\!}$has ${}_{\!}$been ${}_{\!}$pre-trained ${}_{\!}$on ${}_{\!}$the ${}_{\!}$arge-scale ${}_{\!}$pen ${}_{\!}$ime ${}_{\!}$eries ${}_{\!}$rchive ${}_{\!}$(LOTSA) ${}_{\!}$dataset ${}_{\!}$[85], ${}_{\!}$which ${}_{\!}$covers ${}_{\!}$nine ${}_{\!}$varied ${}_{\!}$domains ${}_{\!}$and ${}_{\!}$compiles ${}_{\!}$over ${}_{\!}$27 ${}_{\!}$billion ${}_{\!}$timestamped ${}_{\!}$instances. ${}_{\!}$he ${}_{\!}$spatial ${}_{\!}$decoder ${}_{\!}$first ${}_{\!}$uses ${}_{\!}$a ${}_{\!}$projection ${}_{\!}$block ${}_{\!}$from ${}_{\!}$lama ${}_{\!}$[78, 4] ${}_{\!}$to ${}_{\!}$encode ${}_{\!}$the ${}_{\!}$context ${}_{\!}$description ${}_{\!}$of ${}_{\!}$the ${}_{\!}$input ${}_{\!}$signal, ${}_{\!}$followed ${}_{\!}$by ${}_{\!}$a ${}_{\!}$leaner ${}_{\!}$transformation. ${}_{\!}$utputs ${}_{\!}$are ${}_{\!}$fed ${}_{\!}$into ${}_{\!}$a ${}_{\!}$cross-attention ${}_{\!}$layer, ${}_{\!}$where ${}_{\!}$ey ${}_{\!}$and ${}_{\!}$alue ${}_{\!}$are ${}_{\!}$derived ${}_{\!}$from ${}_{\!}$the ${}_{\!}$contextual ${}_{\!}$embedding ${}_{\!}$and ${}_{\!}$query ${}_{\!}$stems ${}_{\!}$from ${}_{\!}$temporal ${}_{\!}$features, ${}_{\!}$to ${}_{\!}$generate ${}_{\!}$the ${}_{\!}$final ${}_{\!}$spatial ${}_{\!}$features. ${}_{\!}$e concatenate ${}_{\!}$the ${}_{\!}$spatial ${}_{\!}$and ${}_{\!}$temporal ${}_{\!}$features ${}_{\!}$before ${}_{\!}$feeding ${}_{\!}$them ${}_{\!}$into ${}_{\!}$a ${}_{\!}$linear ${}_{\!}$layer ${}_{\!}$to ${}_{\!}$produce ${}_{\!}$the ${}_{\!}$final, ${}_{\!}$augmented ${}_{\!}$input ${}_{\!}$signals.
@@
This section elaborates on a detailed analysis of quantitative results in Table S2, focusing specifically on in-context learning [12] performance and a runtime assessment of the models under investigation. Initially, we present supplementary in-context learning results obtained directly from various LLMs (i.e., Llama 2 [79], Llama 3 [4], and Claude 3 Opus [6]). These findings indicate that, even without an additional fine-tuning process, the LLMs exhibit substantial proficiency in the in-context learning scheme within the ICF task. For instance, Claude 3 Opus [6] achieves CAE scores of 12.19, 10.67, and 9.46 for the 1-shot, 2-shot, and 3-shot scenarios, respectively. It is amply demonstrated that the vanilla LLM has the ability to make inferences and predictions on empirical scientific data even if it is not fine-tuned at all. This underscores that our approach, leveraging these LLMs, represents a notable advancement, particularly in forecasting the energy dynamics of hot electrons.
Furthermore, it is pertinent to emphasize the economic and operational advantages of our computational approach over traditional physical experiments. Specifically, conducting a single ICF experiment typically incurs costs upwards of one million US dollars. Conversely, computational simulations such as a $150\mu m$ PIC simulation [13] require extensive computational resources, amounting to the utilization of 19,584 cores of CPU over a period of 10 hours. In stark contrast, our model necessitates significantly less computational time and resources, requiring only 30 minutes on 2 NVIDIA A100 GPUs for training and only 3~4 seconds for inference with much higher predictive accuracy compared to the PIC simulation. This comparison not only underscores the cost-effectiveness of our approach but also its efficiency and practicality in other scientific applications where computational resource constraints are a critical factor.
Table S1: Quantitative results on Fusion4AI test split for hot electron energy forecasting (see §3.1 for details). Train Time refer to training for the designated task, and Infer Time refer to the amount of time used to predict one case. Note that 3-shot experiments could not be performed on Llama series of models due to the limitation of context window.
@@
PIC Simulation [13] - - $>10$ hrs 2.88 0.20 0.13
LSTM [25] 81.6K ~5 mins $<1$ s 5.82 0.35 0.35
Autoformer [86] 120.4K ~8 mins $<1$ s 5.79 0.35 0.34
GPT4TS [94] 1.5 B ~22 mins ~2 s 3.34 0.18 0.14
Time-LLM [31] 7 B ~20 mins ~3 s 3.48 0.18 0.15
Llama 2 (1-Shot) [79] 7 B - ~3 s 471.22 15.80 14.92
Llama 2 (2-Shot) [79] 7 B - ~5 s 30.33 0.95 0.91
Llama 2 (1-Shot) [79] 70 B - ~7 s 20.63 0.64 0.63
Llama 2 (2-Shot) [79] 70 B - ~8 s 16.42 0.51 0.50
Llama 3 (1-Shot) [4] 8 B - ~4 s 583.14 17.70 16.97
Llama 3 (2-Shot) [4] 8 B - ~6 s 26.66 0.83 0.81
Llama 3 (1-Shot) [4] 70 B - ~14 s 72.35 1.02 1.15
Llama 3 (2-Shot) [4] 70 B - ~19 s 13.62 0.40 0.39
Claude 3 Opus (1-Shot) [6] 137 B - ~12 s 12.19 0.39 0.39
Claude 3 Opus (2-Shot) [6] 137 B - ~17 s 10.67 0.38 0.37
Claude 3 Opus (3-Shot) [6] 137 B - ~20 s 9.46 0.37 0.36
RCRK [17] 106K ~2 mins $<1$ s 4.31 0.28 0.22
HoGRC [43] 394K ~4 mins $<1$ s 4.20 0.25 0.22
NGRC [20] 157K ~2 mins $<1$ s 4.28 0.27 0.23
Fusion-LLM (Llama-2) 7B ~30 mins ~3 s 2.15 0.14 0.12
Fusion-LLM (Llama-3) 8B ~30 mins ~4 s 1.90 0.14 0.11
@@
This section expands to include more qualitative results that help to understand the capabilities and effectiveness of this model. Initially, we release all visualized prediction results of our model Fusion-LLM on test split of our dataset Fusion4AI in Fig. S2. From these qualitative results, it can be found that our model achieves accurate predictions on all unseen data, especially conforming to the temporal and spatial characteristics of the predicted targets, which is crucial for physicists to apply our model as a tool in the design of real-world ICF shots.
@@
Recall that we describe in §S2, the direct prediction of ICF tasks by vanilla LLM use in-context learning method without fine-tuning on our data is not as good as fine-tuned methods in terms of the quantitative value of metric, but it can in fact predict more meaningful patterns than traditional methods such as LSTM [25] and Autoformer [86]. As illustrated in Figure S1, the LLM without fine-tune can infer approximate predictions with reference to the 1 to 3 examples provided, as compared to LSTM [25] and Autoformer [86] which can only predict straight lines close to 0 with patterns. This proves that the vanilla LLMs themselves already contain the capability to infer specific empirical scientific data, which is the core reason we chose LLMs as the reservoir of our Fusion-LLM.
@@
In this section, we examine a significant outlier with the largest error forecasts generated by the Fusion-LLM on the test split. This particular instance serves as a critical case study for understanding the limitations and challenges faced by our model. Fig. S3 illustrates that the shot markedly deviates from the typical scenarios. Notably, this shot exhibits an exceptionally low peak hot electron energy, registering less than 0.15, whereas the majority of other cases yield values ranging between 0.25 and 0.5 under a comparable input laser profile. This anomaly categorizes this shot as an out-of-distribution (OOD) instance. The limited volume of training data available for Fusion-LLM is a plausible explanation for the model’s diminished performance on this OOD data. In scenarios where training data is sparse, the model’s capability to generalize to new, especially atypical, data points is inherently restricted. Consequently, this case highlights the importance of enhancing the dataset’s diversity and volume for ICF tasks. We hope the community can share more data points to improve and enlarge Fusion4AI dataset (§3.1) together.

### v3

@@ 3.2 Diagnostic Experiment
# of samples Ours HoGRC NGRC Time-LLM
@@
# of epochs Ours HoGRC NGRC Time-LLM
@@
Reservoir and LLM Comparison. To thoroughly explore the training effectiveness of our methodology under conditions of limited sample availability, we perform a comparative analysis using a variable number of training samples with two concurrent reservoir methods (Li et al., 2024; Gauthier et al., 2021) and one LLM-based method (Jin et al., 2023) in Table 2d using CAE. The empirical findings demonstrate that our approach consistently outperforms all competing strategies across various sample configurations. Notably, this superior performance and training effectiveness are evident even with as few as 20 samples. Such robust efficacy is critical in scientific AI applications, where datasets are often constrained in size.
In addition, we study the training efficiency of our approach in contrast to the above strong baselines (Li et al., 2024; Gauthier et al., 2021; Jin et al., 2023) in Table 2e, across various training epochs. The experimental outcomes illustrate that our approach not only outperforms its counterparts but also demonstrates superior efficiency. Specifically, our method is capable of achieving comparable or superior performance in a significantly reduced training duration. For instance, our model requires only 10 epochs to achieve better performance than other methods that require 100 epochs. This enhanced efficiency in training is particularly significant, as it demonstrates the potential of our approach to deliver robust performance swiftly, thereby facilitating more expedient research in practical scenarios.
@@ 4 Related Work and Discussion
AI for Science. AI has increasingly become a vital tool in advancing scientific discovery, playing a central role in recent breakthroughs across various fields (Jumper et al., 2021; LeCun et al., 2015; Reichstein et al., 2019; Ali et al., 2024). The trajectory of AI’s involvement in scientific research began with elementary data analysis techniques, such as rule-based systems (Breiman, 2001; Safavian & Landgrebe, 1991), Bayesian methods (Frank et al., 2000), analogy-based approaches (Hearst et al., 1998; Jain et al., 1999; Tenenbaum et al., 2000), evolutionary algorithms (Kennedy & Eberhart, 1995; Dietterich, 2000), and connectionist models (Weisberg, 2005; LeCun et al., 2015). These methods laid the foundation for AI’s contributions to scientific exploration, and have since evolved into sophisticated models, including deep learning (He et al., 2016), transformers (Vaswani et al., 2017), and foundation models (Dosovitskiy et al., 2021; Bommasani et al., 2021). These advancements have empowered scientists to tackle complex problems with greater accuracy and efficiency.
Distinct from traditional AI approaches in scientific exploration, our work represents a one of the first efforts to empower a crucial scientific domain in ICF with LLMs, which has previously relied on traditional computational techniques. Despite lacking explicit first-principle rules, our LLM-based models exhibit remarkable predictive accuracy on real-world data, offering a promising alternative to traditional simulations and costly experimental data collection. Specifically, our method demonstrates the potential to revolutionize processes such as ICF, by providing reliable, data-driven insights that guide experimental setups with reduced reliance on conventional computation methods.
@@
Plasma Physics for Fusion. LPI is an important area of study within plasma physics for fusion due to its potential to decrease the efficiency of the implosion process. Stimulated Raman scattering (SRS) and stimulated Brillouin scattering (SBS) (Kirkwood et al., 1999) can cause the reflection of laser beams. Cross-beam energy transfer (CBET) can adversely influence the symmetry of implosion (Igumenshchev et al., 2010). Two-plasmon decay (TPD) can effectively generate hot electrons, which can preheat the target, increase the shell entropy, and diminish the implosion efficiency. (Smalyuk et al., 2008a; Goncharov et al., 2008; Craxton et al., 2015; Radha et al., 2016). In shock ignitions (Betti et al., 2007; Perkins et al., 2009), hot electrons can also deposit their energy in the compressed shell, thereby enhancing the ignition shock and aiding ignition processes. Understanding LPI physics and establishing predictive models for hot electron generation in direct drive ICF are crucial. In response to these issues and challenges, extensive experiments and simulations (Smalyuk et al., 2008a; Goncharov et al., 2008; Craxton et al., 2015; Radha et al., 2016; Betti et al., 2007; Perkins et al., 2009) are necessary, which are both costly and time-consuming.
To address these challenges, our AI-based approach offers a promising alternative. By implementing a streamlined LLM pipeline, our model acts as a Computational Reservoir for time-series forecasting, capturing domain-specific knowledge and generalizing within the ICF task. This enables LLMs to assist in ICF design, reducing reliance on costly experiments and simulations. Empirical results (see §3) suggest that LLMs could revolutionize predictive modeling in plasma physics, providing rapid, cost-effective insights based on generalizable knowledge.
Notably, our LPI-LLM for predicting hot electron-induced hard X-rays would provide a useful framework for predicting other experimental data, such as neutron yields, with different but equally intricate underlying physics. Through these applications, LPI-LLM has the potential to become ICF-LLM, significantly advancing fusion research, paving the way for new insights and advancements in sustainable energy production.
Reservoir Computing. Traditional machine learning techniques (Sutton, 1988; Arel et al., 2010; Ahmed et al., 2010; Williams & Rasmussen, 2006; Samuel, 1959; Sapankevych & Sankar, 2009; Kadous, 1999) for scientific data often rely on transforming temporal inputs into high-dimensional state spaces using nonlinear mappings. RC introduced through key advancements like Echo State Networks (Jaeger, 2007; Lukoševičius, 2012) and Liquid State Machines (Maass, 2011; Zhang et al., 2015), offers a compelling alternative to these traditional methods. RC operates by employing a fixed, untrained dynamic system—the reservoir—which processes input signals into rich representations. A notable advantage of RC is that only the readout layer is trained, drastically reducing computational overhead and simplifying the learning process. This efficiency makes RC particularly suited for large-scale scientific applications, where the volume and complexity of data demand scalable and low-cost solutions (Nakajima & Fischer, 2021; Schrauwen et al., 2007; Gauthier et al., 2021; Lukoševičius & Jaeger, 2009).
Our study proposes a paradigm shift by integrating LLMs into the RC framework, advancing beyond the limitations of traditional RNN-based reservoirs (Gauthier et al., 2021; Lukoševičius & Jaeger, 2009). Unlike standard reservoirs that may struggle with highly dynamic or noisy data, our method leverages the pre-trained knowledge and adaptive reasoning of LLMs, significantly enhancing RC’s capacity to handle complex scientific datasets like LPI. This integration not only improves the system’s ability to capture intricate temporal patterns but also minimizes the need for extensive parameter tuning. As a result, our approach offers a more powerful, adaptable, and efficient solution for scientific tasks, setting a new benchmark for RC-based methodologies.
Time-series Forecasting. In various scientific fields, time-series data plays a pivotal role, including ICF, where the temporal evolution of laser intensity significantly impacts target behavior and hot electron emission. Traditional methods for time-series analysis (Sutton, 1988; Arel et al., 2010; Williams & Rasmussen, 2006; Samuel, 1959; Sapankevych & Sankar, 2009; Kadous, 1999) typically involve the transformation of temporal inputs into high-dimensional spaces. However, these methods frequently encounter limitations when dealing with highly dynamic and complex data (Ahmed et al., 2010). Recently, transformer-based models for time-series forecasting, such as Temporal Fusion Transformers (TFT) (Lim et al., 2021), Informer (Zhou et al., 2021) and Autoformer (Wu et al., 2021), have demonstrated significant improvements by effectively capturing long-term dependencies and scaling to large temporal datasets. These models employ attention mechanisms that enhance accuracy, particularly in forecasting long-horizon data sequences. With the rise of large-scale models, the use of LLMs for time-series forecasting has emerged as a promising approach. For instance, Time-LLM (Jin et al., 2023) and GPT4TS (Zhou et al., 2023) have leveraged the vast pre-trained knowledge bases of LLMs to model complex temporal patterns more efficiently than previous methods. These LLM-based approaches benefit from their ability to handle varied, intricate time-series data with minimal task-specific fine-tuning, offering a flexible solution that adapts to diverse forecasting scenarios.
Conceptually different than these prior arts, our method introduces a novel architecture that synergistically combines multiple components to address the unique challenges of LPI forecasting in ICF. The Fusion-Specific Prompts strategically connect the LLM’s vast knowledge base to ICF-specific nuances, enhancing the model’s ability to interpret plasma physics phenomena. Our Signal-Digesting Channels, comprising temporal and spatial encoders, are specifically designed to capture the complex temporal patterns and critical landscape features of laser intensity signals in ICF. This multi-faceted approach allows LPI-LLM to more precisely model the intricate dynamics and uncertainties inherent in LPI data. By integrating these components, our framework achieves superior adaptability and robustness in time-series forecasting for ICF applications, overcoming limitations of previous methods in handling high-dimensional, volatile data scenarios typical in plasma physics.
@@ 5 Conclusion
Fusion energy stands as a pivotal pathway toward advancing human civilization to a Type I status on the Kardashev scale (Kardashev, 1964). The key to realizing this potential lies in mastering Inertial Confinement Fusion, where understanding laser-plasma instabilities is paramount. To address this challenge, we present LPI-LLM, a groundbreaking framework merging LLMs with reservoir computing. Our approach not only provides a cost-effective solution but also emerges as a top-tier contender in forecasting hot electron dynamics, offering invaluable insights for plasma scientists in ${}_{\!}$refining ${}_{\!}$ ${}_{\!}$designs. ${}_{\!}$eyond ${}_{\!}$its ${}_{\!}$immediate ${}_{\!}$impact ${}_{\!}$on ${}_{\!}$, ${}_{\!}$employing ${}_{\!}$s for ${}_{\!}$scientific ${}_{\!}$exploration ${}_{\!}$holds ${}_{\!}$promise ${}_{\!}$for ${}_{\!}$${}_{\!}$cross-domain ${}_{\!}$applications, ${}_{\!}$potentially ${}_{\!}$catalyzing ${}_{\!}$advancements ${}_{\!}$in ${}_{\!}$-driven ${}_{\!}$scientific ${}_{\!}$endeavors.
6 Ethics
7 Reproducibility
LPI-LLM is implemented in PyTorch (Paszke et al., 2019). Experiments are conducted on two NVIDIA A100-40GB GPUs. To guarantee reproducibility, we fully describe our approach in §2 and implementation details in Appendix §S1. Our full implementation of code, model weights, and test split of the dataset are also submitted with this paper for reproduction.
@@ 7 Reproducibility
The overall pipeline of LPI-LLM is shown in Fig. 3. Experiments are conducted on two NVIDIA A100-40GB GPUs. For our approach, we keep all parameters of the LLMs and most of the SDC frozen during the fine-tuning. Only parameters pertaining to the Prediction Head and partial Spatial Encoder are trainable. The codes and dataset shall be publicly released upon paper acceptance.
  • LPI-LLM is built from Llama 2 7B (Touvron et al., 2023a), Llama 3 8B (AI, 2024), Gemma 2 9B (Team, 2024) and OLMo (Groeneveld et al., 2024) to construct reservoir without tuning.
@@
  • Signal-digesting channels are composed of two components, temporal encoder and spatial encoder. The former one, which incorporates 24 Transformer layers and a linear layer, captures temporal features over the input laser signal. This module has been pre-trained on the Large-scale Open Time Series Archive (LOTSA) dataset (Woo et al., 2024), which covers nine varied domains and compiles over 27 billion timestamped instances. The spatial decoder first uses a projection block from LLMs to encode the context description of the input signal, followed by a leaner transformation. Outputs are fed into a cross-attention layer, where Key and Value are derived from the contextual embedding and query stems from temporal features, to generate the final spatial features. We concatenate the spatial and temporal features before feeding them into a linear layer to produce the final, augmented input signals.
@@
This section elaborates on a detailed analysis of quantitative results in Table S2, focusing specifically on in-context learning (Brown et al., 2020) performance and a runtime assessment of the models under investigation. Initially, we present supplementary in-context learning results obtained directly from various LLMs (i.e., Llama 2 (Touvron et al., 2023b), Llama 3 (AI, 2024), and Claude 3 Opus (Anthropic, 2024)). These findings indicate that, even without an additional fine-tuning process, the LLMs exhibit substantial proficiency in the in-context learning scheme within the ICF task. For instance, Claude 3 Opus (Anthropic, 2024) achieves CAE scores of 12.19, 10.67, and 9.46 for the 1-shot, 2-shot, and 3-shot scenarios, respectively. It is amply demonstrated that the vanilla LLM has the ability to make inferences and predictions on empirical scientific data even if it is not fine-tuned at all. This underscores that our approach, leveraging these LLMs, represents a notable advancement, particularly in forecasting the energy dynamics of hot electrons.
Furthermore, it is pertinent to emphasize the economic and operational advantages of our computational approach over traditional physical experiments. Specifically, conducting a single ICF experiment typically incurs costs upwards of one million US dollars. Conversely, computational simulations such as a $150\mu m$ PIC simulation (Cao et al., 2022) require extensive computational resources, amounting to the utilization of 19,584 cores of CPU over a period of 10 hours. In stark contrast, our model necessitates significantly less computational time and resources, requiring only 30 minutes on 2 NVIDIA A100 GPUs for training and only 3$\sim$4 seconds for inference with much higher predictive accuracy compared to the PIC simulation. This comparison not only underscores the cost-effectiveness of our approach but also its efficiency and practicality in other scientific applications where computational resource constraints are a critical factor.
Table S1: Quantitative results on LPI4AI test split for hot electron energy forecasting (see §3.1 for details). Train Time refers to training for the designated task, and Infer Time refers to the amount of time used to predict one case. Note that 3-shot experiments could not be performed on Llama series of models due to the limitation of the context window.
@@
PIC Simulation - - $>10$ hrs 2.88 0.20 0.13
LSTM 81.6K $\sim$5 mins $<1$ s 5.82 0.35 0.35
Autoformer 120.4K $\sim$8 mins $<1$ s 5.79 0.35 0.34
GPT4TS 1.5 B $\sim$22 mins $\sim$2 s 3.34 0.18 0.14
Time-LLM 7 B $\sim$20 mins $\sim$3 s 3.48 0.18 0.15
Llama 2 (1-Shot) 7 B - $\sim$3 s 471.22 15.80 14.92
Llama 2 (2-Shot) 7 B - $\sim$5 s 30.33 0.95 0.91
Llama 2 (1-Shot) 70 B - $\sim$7 s 20.63 0.64 0.63
Llama 2 (2-Shot) 70 B - $\sim$8 s 16.42 0.51 0.50
Llama 3 (1-Shot) 8 B - $\sim$4 s 583.14 17.70 16.97
Llama 3 (2-Shot) 8 B - $\sim$6 s 26.66 0.83 0.81
Llama 3 (1-Shot) 70 B - $\sim$14 s 72.35 1.02 1.15
Llama 3 (2-Shot) 70 B - $\sim$19 s 13.62 0.40 0.39
Claude 3 Opus (1-Shot) 137 B - $\sim$12 s 12.19 0.39 0.39
Claude 3 Opus (2-Shot) 137 B - $\sim$17 s 10.67 0.38 0.37
Claude 3 Opus (3-Shot) 137 B - $\sim$20 s 9.46 0.37 0.36
RCRK 106 K $\sim$2 mins $<1$ s 4.31 0.28 0.22
HoGRC 394 K $\sim$4 mins $<1$ s 4.20 0.25 0.22
NGRC 157 K $\sim$2 mins $<1$ s 4.28 0.27 0.23
LPI-LLM (Gemma-2) 9 B $\sim$30 mins $\sim$4 s 2.04 0.14 0.12
LPI-LLM (OLMo) 7 B $\sim$30 mins $\sim$3 s 1.97 0.14 0.12
LPI-LLM (Llama-2) 7 B $\sim$30 mins $\sim$3 s 2.15 0.14 0.12
LPI-LLM (Llama-3) 8 B $\sim$30 mins $\sim$4 s 1.90 0.14 0.11
@@
This section expands to include more qualitative results that help to understand the capabilities and effectiveness of this model. Initially, we release all visualized prediction results of our model LPI-LLM on test split of our dataset LPI4AI in Fig. S2. From these qualitative results, it can be found that our model achieves accurate predictions on all unseen data, especially conforming to the temporal and spatial characteristics of the predicted targets, which is crucial for physicists to apply our model as a tool in the design of real-world ICF shots.
@@
Recall that we describe in §S2, the direct prediction of ICF tasks by vanilla LLM use in-context learning method without fine-tuning on our data is not as good as fine-tuned methods in terms of the quantitative value of metric, but it can in fact predict more meaningful patterns than traditional methods such as LSTM (Hochreiter & Schmidhuber, 1997) and Autoformer (Wu et al., 2021). As illustrated in Figure S1, the LLM without fine-tune can infer approximate predictions with reference to the 1 to 3 examples provided, as compared to LSTM (Hochreiter & Schmidhuber, 1997) and Autoformer (Wu et al., 2021) which can only predict straight lines close to 0 with patterns. This proves that the vanilla LLMs themselves already contain the capability to infer specific empirical scientific data, which is the core reason we chose LLMs as the reservoir of our LPI-LLM.
@@
In this section, we examine a significant outlier with the largest error forecasts generated by the LPI-LLM on the test split. This particular instance serves as a critical case study for understanding the limitations and challenges faced by our model. Fig. S3 illustrates that the shot markedly deviates from the typical scenarios. Notably, this shot exhibits an exceptionally low peak hot electron energy, registering less than 0.15, whereas the majority of other cases yield values ranging between 0.25 and 0.5 under a comparable input laser profile. This anomaly categorizes this shot as an out-of-distribution (OOD) instance. The limited volume of training data available for LPI-LLM is a plausible explanation for the model’s diminished performance on this OOD data. In scenarios where training data is sparse, the model’s capability to generalize to new, especially atypical, data points is inherently restricted. Consequently, this case highlights the importance of enhancing the dataset’s diversity and volume for ICF tasks. We hope the community can share more data points to improve and enlarge LPI4AI dataset (§3.1) together.